Circular No. 6903

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bw IN ESO 184-G82 S. R. Kulkarni and J. S. Bloom, California Institute of Technology;

D. A. Frail, National Radio Astronomy Observatory; and R. Ekers, M. Wieringa, R. Wark, and J. L. Higdon, Australian Telescope National Facility, report: “Within the localization of GRB 980425 (IAUC 6884), Galama et al. (IAUC 6895) reported a possible supernova candidate for which Wieringa et al. (IAUC 6896) saw a brightening radio source. The object appears to be a young supernova, possibly of type-Ib (IAUC 6901), which exploded on or around Apr. 24, the epoch of GRB 980425. For an assumed expansion speed of 20 000 km/s and a distance of 44 Mpc to the host galaxy of the supernova (from the redshift given on IAUC 6896), we derive a brightness temperature of 3 × 10^{14} K from the observed 39 mJy at 6 cm on May 5 (IAUC 6896). This is in excess of the usual Compton limit of 10^{12} K. Despite this, no x-ray emission is seen (GCN 69). Thus we are forced to invoke a relativistic expansion speed that results in a larger source size and correspondingly smaller brightness temperature. We suggest that the radio emission arises in a relativistic shock and the optical emission in a standard low-velocity shock. We urge observers to carry out higher-frequency radio observations and infrared observations, as the radio spectrum may rapidly evolve. The model predicts that the radio source should not exhibit diffractive scintillation. VLBI observations would be of great value.”

Visual magnitude estimates by B. Monard, Pretoria, South Africa: May 10.14 UT, 13.5; 13.15, 13.5.

SUPERNOVA 1988ae IN MCG +02-53-002 J. Mueller reports her discovery of an apparent supernova (mag ∼ 16.5)

on a plate taken with the 1.2-m Oschin Schmidt Telescope on 1988 Sept. 15 in the course of the second Palomar Sky Survey. SN 1988ae is located at α = 20^{h}44^{m}09^{s}.0, δ = +12^{o}24′53′′ (equinox 2000.0), which is 10′′ west and 15′′ south of the center of MCG +02-53-002. The star also appears on a Sky Survey plate taken by Mueller on 1988 Aug. 12 at mag ∼ 17. Sky Survey plates taken on 1988 Aug. 6 and 1990 Aug. 18 do not show SN 1988ae, neither does the object appear on the Digital Sky Survey.

1998 May 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green